OpenStudy (anonymous):

Quadatric equation is given by (2a+1)x^2-ax+a-1=0, and two real roots are p and q. Find the value range of a such that p >1 and q <1.

OpenStudy (anonymous):

Do you have the answers? If so, could you show me to see I got the same answer.

OpenStudy (anonymous):

@Azteck no i don't know how to solve this :/

OpenStudy (anonymous):

\[ax^2-bx+c=0\] \[\alpha + \beta=\frac{ -b }{ a }\] \[\alpha \beta=\frac{ c }{ a }\] ---------------------------------------------------------- \[p+q=\frac{-(-a)}{2a+1}\] \[pq=\frac{ a-1 }{ 2a+1 }\] DISCRIMINANT \[b^2-4ac \ge 0\] ^real roots means greater than or equal to zero. \[(-a)^2-4(2a+1)(a-1) \ge 0\] \[a^2-4(2a^2-a-1) \ge 0\] \[a^2-8a^2+4a+4 \ge 0\] \[-7a^2+4a+4 \ge 0\] \[7a^2-4a-4 \le 0\]

OpenStudy (anonymous):

I think you need to use the discriminant plus the sum of roots.

OpenStudy (anonymous):

and product of roots.

OpenStudy (anonymous):

@Azteck Thank u so much for ur answer. It's really helpful:)

OpenStudy (anonymous):

But at one point I'm still in doubt. How can I make sure the two roots are satisfying this condition: p >1 and q <1

OpenStudy (anonymous):

I'm not entirely sure. I'm trying to think.

OpenStudy (anonymous):

Do you have the correct answers. Do you have an answer sheet or something like that?

OpenStudy (anonymous):

Nope, unfortunately I have only the questions sheets :/

OpenStudy (anonymous):

I don't know if the p and q as roots are just extraneous information. I think I need assistance. @satellite73 Do you know how to answer this question?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (shubhamsrg):

firstly, let us go for real roots, as @azteck rightly did, we have 7a^2 - 4a - 4 <= 0 Note we don't want equal roots, hence 7a^2 - 4a - 4 < 0 Let us try to say from the graph, If, 2a+1 >0, the graph will be an upward parabola like this :  If 2a-1 <0, the graph will have be a downward parabola, like this:  Where we have used the info that 1 lies between both roots, from either of the graphs, you can easily see that our 2nd required condition would be (2a+1)(f(1)) <0 => (2a+1)(2a + 1 - a +a -1 )<0 =>(2a+1)(a) <0 Solve this inequality, and the previous one, and take the common points as the final answer. Hope that helps.